OpenStudy (anonymous):

how do u find( A^-1)^-1 or the inverse of A inverse

OpenStudy (anonymous):

its a 3x3

OpenStudy (anonymous):

A

OpenStudy (asnaseer):

@jimmyrep is correct - the inverse of an inverse just gets you back to the original.

OpenStudy (asnaseer):

think of it using fractions, the inverse of 3 is 1/3. the inverse of 1/3 is:\[\frac{1}{\frac{1}{3}}=1\div\frac{1}{3}=1\times\frac{3}{1}=3\]

OpenStudy (anonymous):

i knw it gets you back to the original but do you go through the whole 1/det of a-1 by its adjoint or do u just transpose it

OpenStudy (asnaseer):

if you know how to get \(A^{-1}\) from \(A\), then use exactly the same method on \(A^{-1}\) to invert it and you will find you get back to \(A\).

OpenStudy (anonymous):

oh thats waht i was asking...damn thats gonna take long...also to find determinant of matrices u use diff formulas u can use cofactors, ero's i dont understand the one that seems as tho its cofactor but its not could you explain

OpenStudy (anonymous):

eg a11a22a33-a11a32a23-a12a21a33+a12a31a23+a13a21a32-a13a31a22 that...my teacher didnt quite get through to me with this one when do u use this

OpenStudy (asnaseer):

see here for an example of how to invert a 3x3 matrix: http://www.wikihow.com/Inverse-a-3X3-Matrix

OpenStudy (asnaseer):

this site shows you the inverse using the notation that you were using: http://www.dr-lex.be/random/matrix_inv.html